# Metric Distortion of Line-up Elections: The Right Person for the Right Job

### v1

@@ 4 Knowing Ordinal Position Preferences
There exists a mechanism that achieves a distortion of $3$ for ($k$,$1$)-line-up elections while knowing only the position preferences. Moreover, no deterministic mechanism can have smaller worst-case distortion while knowing only the position preferences, even on a line and with only 2 candidates.
@@
After considering the case where we have $k$ candidates and one position, we now extend the setting to $k$ candidates and $k$ positions. In this case, each candidate will be matched to one position and vice versa. We will first show that serial dictatorship gives a trade off between the cost of the position side of the matching and the cost of the voter side of the matching. To do that we will revisit Lemma 3.4 and derive a new lemma.
@@
Serial dictatorship gives a distortion of at most $3-\frac{1}{2^{k-2}}$ for ($k$,$k$)-line-up elections.
@@
Let $M_{1}$ be the matching given by serial dictatorship and $M^{*}$ be the optimal matching. We know from Theorem 1 of Caragiannis et al. [2022a] that serial dictatorship for bipartite perfect matching gives a distortion of at most $12^{k}-$ 1 so we have $\operatorname{cost}_{P}(M_{1})\leq(2^{k}-1)\operatorname{cost}_{P}(M^{*})$. Therefore, due to the Lemma above, and the fact that for ($k$,$k$)-line-up elections we have $\operatorname{cost}_{V}(M_{1})=\operatorname{cost}_{V}(M^{*})$, our distortion is given by
$\min\left\{\frac{3\operatorname{cost}_{V}(M^{*})+\operatorname{cost}_{P}(M^{*})}{\operatorname{cost}_{V}(M^{*})+\operatorname{cost}_{P}(M^{*})},\frac{\operatorname{cost}_{V}(M^{*})+(2^{k}-1)\operatorname{cost}_{P}(M^{*})}{\operatorname{cost}_{V}(M^{*})+\operatorname{cost}_{P}(M^{*})}\right\}.$
@@
Thus $\frac{\operatorname{cost}_{P}(M_{1})+\operatorname{cost}_{V}(M_{1})}{\operatorname{cost}_{P}(M^{*})+\operatorname{cost}_{V}(M^{*})}\leq\frac{\operatorname{cost}_{P}(M^{*})+3(2^{k-1}-1)\operatorname{cost}_{P}(M^{*})}{\operatorname{cost}_{P}(M^{*})+(2^{k-1}-1)\operatorname{cost}_{P}(M^{*})}=\frac{3\cdot 2^{k-1}-2}{2^{k-1}}=3-\frac{1}{2^{k-2}}$. ∎
Theorem 4.3 shows that simply by using serial dictatorship we can get a distortion better than 3. For $k=2$ we get a distortion of $2$, and for $k=3$ we get $\frac{5}{2}$, so for small $k$ the distortion remains relatively low despite approaching 3 as $k$ becomes larger.
We now consider the general case where we have $k$ candidates and $\ell$ positions, with $k\geq\ell$. We will now show that simply running serial dictatorship on the set of positions results in an upper bound on the distortion of $5$. Having the extra information on the ordering of candidates by fitness for each position not only allows us to improve the bounds from Theorem 3.10 when we only know voter preferences, but also does so without needing voter preferences at all. Note that serial dictatorship then is also strategy proof, since we do not consider the preferences of the candidates or the voters, and even the positions are not incentivized to lie (if such a thing were possible) to achieve a closer candidate.
@@
Serial dictatorship achieves a distortion of at most $5$ for ($k$$\ell$)-line-up elections.
@@ 5 Knowing Candidate and Position Locations
There exists a deterministic mechanism that gives a distortion of at most $3$ for the ($k$$\ell$)-line-up election while knowing only candidate-position distance pairs. Moreover, no deterministic mechanism can achieve a distortion bound less than $3$ while knowing only position and candidate locations.
@@
To prove the lower bound, consider two instances both with $\mathcal{P}=\{1\}$, $\mathcal{C}=\{A,B\}$ and $d(A,1)=1-\epsilon$, $d(B,1)=1+\epsilon$ and $d(A,B)=2$. One instance will have $n$ voters on $A$ and the other will have $n$ on top of $B$. If we chose $A$ then the second instance will have a distortion of $\lim_{\epsilon\to 0}\frac{n(d(A,B)+d(A,1))}{nd(B,1)}=\lim_{\epsilon\to 0}\frac{3-\epsilon}{1+\epsilon}=3$. If we pick $B$ then the first instance will have a distortion of $\lim_{\epsilon\to 0}\frac{n(d(A,B)+d(B,1))}{nd(A,1)}=\lim_{\epsilon\to 0}\frac{3+\epsilon}{1-\epsilon}=3$. ∎
Much like the mechanisms in Theorem 4.1 and 4.3 this mechanism is strategyproof and does not rely on the preferences of the voters. This means that we can achieve low distortion without surveying voters, and instead focusing on increasing our knowledge of the pairwise distances between candidates and positions. Note that for ($k$,$k$)-line-up elections, the above mechanism trivially computes the optimum solution, since all candidates must always be chosen in such elections, and thus $\operatorname{cost}_{V}(M)$ of any matching $M$ is always the same, and only $\operatorname{cost}_{P}(M)$ needs to be minimized.
@@
There is no deterministic mechanism for the ($k$,$1$)-line-up election problem knowing candidate position distance pairs and with voter preferences that achieves a better distortion better than $\frac{5}{3}$, even on a line and with only $2$ candidates.
@@
Assume there are $k$ candidates. We start by constructing a graph $G=(V,A)$ similar to a weighted majority graph. Instead we orient arcs such that instead of arcs going from winners according to a plurality election, they are oriented based on the winner given by Theorem 5.3. So if $(x,y)\in A$ then $x$ was selected by the mechanism used in Theorem 5.3. First we introduce the following lemma regarding the graph previously described. We then select the client with the largest out degree.
@@
There exists a mechanism that achieves a distortion of $\frac{25}{9}$ for ($k$,$1$)-line-up election.

### v2

@@ 4 Knowing Ordinal Position Preferences
There exists a mechanism that achieves a distortion of $3$ for ($m$,$1$)-line-up elections while knowing only the position preferences. Moreover, no deterministic mechanism can have smaller worst-case distortion while knowing only the position preferences, even on a line and with only 2 candidates.
@@
After considering the case where we have $m$ candidates and one position, we now extend the setting to $m$ candidates and $m$ positions. In this case, each candidate will be matched to one position and vice versa. We will first show that serial dictatorship gives a trade off between the cost of the position side of the matching and the cost of the voter side of the matching. To do that we will revisit Lemma 3.4 and derive a new lemma.
@@
Serial dictatorship gives a distortion of at most $3-\frac{1}{2^{m-2}}$ for ($m$,$m$)-line-up elections.
@@
Let $M_{1}$ be the matching given by serial dictatorship and $M^{*}$ be the optimal matching. We know from Theorem 1 of Caragiannis et al. [2022a] that serial dictatorship for bipartite perfect matching gives a distortion of at most $12^{m}-$ 1 so we have $\operatorname{cost}_{P}(M_{1})\leq(2^{m}-1)\operatorname{cost}_{P}(M^{*})$. Therefore, due to the Lemma above, and the fact that for ($m$,$m$)-line-up elections we have $\operatorname{cost}_{V}(M_{1})=\operatorname{cost}_{V}(M^{*})$, our distortion is given by
$\min\left\{\frac{3\operatorname{cost}_{V}(M^{*})+\operatorname{cost}_{P}(M^{*})}{\operatorname{cost}_{V}(M^{*})+\operatorname{cost}_{P}(M^{*})},\frac{\operatorname{cost}_{V}(M^{*})+(2^{m}-1)\operatorname{cost}_{P}(M^{*})}{\operatorname{cost}_{V}(M^{*})+\operatorname{cost}_{P}(M^{*})}\right\}.$
@@
Thus $\frac{\operatorname{cost}_{P}(M_{1})+\operatorname{cost}_{V}(M_{1})}{\operatorname{cost}_{P}(M^{*})+\operatorname{cost}_{V}(M^{*})}\leq\frac{\operatorname{cost}_{P}(M^{*})+3(2^{m-1}-1)\operatorname{cost}_{P}(M^{*})}{\operatorname{cost}_{P}(M^{*})+(2^{m-1}-1)\operatorname{cost}_{P}(M^{*})}=\frac{3\cdot 2^{m-1}-2}{2^{m-1}}=3-\frac{1}{2^{m-2}}$. ∎
Theorem 4.3 shows that simply by using serial dictatorship we can get a distortion better than 3. For $m=2$ we get a distortion of $2$, and for $m=3$ we get $\frac{5}{2}$, so for small $m$ the distortion remains relatively low despite approaching 3 as $m$ becomes larger.
We now consider the general case where we have $m$ candidates and $\ell$ positions, with $m\geq\ell$. We will now show that simply running serial dictatorship on the set of positions results in an upper bound on the distortion of $5$. Having the extra information on the ordering of candidates by fitness for each position not only allows us to improve the bounds from Theorem 3.10 when we only know voter preferences, but also does so without needing voter preferences at all. Note that serial dictatorship then is also strategy proof, since we do not consider the preferences of the candidates or the voters, and even the positions are not incentivized to lie (if such a thing were possible) to achieve a closer candidate.
@@
Serial dictatorship achieves a distortion of at most $5$ for ($m$$\ell$)-line-up elections.
@@ 5 Knowing Candidate and Position Locations
There exists a deterministic mechanism that gives a distortion of at most $3$ for the ($m$$\ell$)-line-up election while knowing only candidate-position distance pairs. Moreover, no deterministic mechanism can achieve a distortion bound less than $3$ while knowing only position and candidate locations.
@@
To prove the lower bound, consider two instances both with $\mathcal{P}=\{1\}$, $\mathcal{C}=\{A,B\}$ and $d(A,1)=1-\epsilon$, $d(B,1)=1+\epsilon$ and $d(A,B)=2$. One instance will have $n$ voters on $A$ and the other will have $n$ on top of $B$. If we choose $A$, then the second instance will have a distortion of $\lim_{\epsilon\to 0}\frac{n(d(A,B)+d(A,1))}{nd(B,1)}=\lim_{\epsilon\to 0}\frac{3-\epsilon}{1+\epsilon}=3$. If we choose $B$ then the first instance will have a distortion of $\lim_{\epsilon\to 0}\frac{n(d(A,B)+d(B,1))}{nd(A,1)}=\lim_{\epsilon\to 0}\frac{3+\epsilon}{1-\epsilon}=3$. ∎
Much like the mechanisms in Theorem 4.1 and 4.3 this mechanism is strategyproof and does not rely on the preferences of the voters. This means that we can achieve low distortion without surveying voters, and instead focusing on increasing our knowledge of the pairwise distances between candidates and positions. Note that for ($m$,$m$)-line-up elections, the above mechanism trivially computes the optimum solution, since all candidates must always be chosen in such elections, and thus $\operatorname{cost}_{V}(M)$ of any matching $M$ is always the same, and only $\operatorname{cost}_{P}(M)$ needs to be minimized.
@@
There is no deterministic mechanism for the ($m$,$1$)-line-up election problem knowing candidate position distance pairs and with voter preferences that achieves a better distortion better than $\frac{5}{3}$, even on a line and with only $2$ candidates.
@@
Assume there are $m$ candidates. We start by constructing a graph $G=(V,A)$ similar to a weighted majority graph. Instead we orient arcs such that instead of arcs going from winners according to a plurality election, they are oriented based on the winner given by Theorem 5.3. So if $(x,y)\in A$ then $x$ was selected by the mechanism used in Theorem 5.3. First we introduce the following lemma regarding the graph previously described. We then select the client with the largest out degree.
@@
There exists a mechanism that achieves a distortion of $\frac{25}{9}$ for ($m$,$1$)-line-up election.